The 1998-98 NSERC

Subatomic Physics Competition

Pekka K. Sinervo University of Toronto

1. Grant Selection Committee

2. Background from Past Years

3. Site Reviews

4. Long Range Planning

5. NSERC Reallocation Steering Committee

TRIUMF Users Meeting 3 December 1997

Grant Selection Committee

The committee membership is

• Pekka Sinervo (chair)

• Nathan Wiess (UBC)

• Pat Kalnyiak (Carleton)

• Jean-Eudes Augustin (IN2P3)

• Sam Austin (Michigan)

• Walter Henning (Argonne)

• Barry Holstein (Massachusetts)

• Vera Luth (SLAC)

• Hay Boon Mak (Queens)

• Richard Milner (MIT)

• Doug Stairs (McGill)

• Michel Vetterli (TRIUMF)

Background from Previous Years

Last year

• Based on Fenton Plan (5-year timeframe)

• Developed a five year plan – Capital spending for SNO, ISAC and ATLAS

– Operating support for ongoing and emerging programme

Now proceeding in this five year framework

• First year where GSC will be developing 5-year model

• Meant to supplement and not replace community based planning process